OpenStudy (anonymous):

Which value is a solution for the equation tan x/2=0 3pi/2 pi 2pi pi/2

OpenStudy (mathstudent55):

Where is the tangent equal to zero?

OpenStudy (anonymous):

I'm not sure?

OpenStudy (mathstudent55):

Think of the tangent as being \(\tan x = \dfrac{\sin x}{\cos x}\) The tangent is zero where the sine is zero. At what values is the sine zero?

OpenStudy (anonymous):

would it also be zero?

OpenStudy (mathstudent55):

The function y = sin x is zero at all integer multiples pf pi.